OpenStudy (anonymous):

log x √x^2+1 i have to write as a sum, i don't know where to start

OpenStudy (anonymous):

this is what it looks like \[\log x \sqrt{x^{2}+1}\]

ganeshie8 (ganeshie8):

log(ab) = log(a) + log(b) can you start now ?

OpenStudy (anonymous):

so it s be log x + log √x^2+1?

ganeshie8 (ganeshie8):

thats not complete yet. you can expand second term further. \[\log m ^{n} = n \log m\]

OpenStudy (anonymous):

Hints: \(\Large \log(\color{red}a \color{blue}b) = \log (\color{red} a ) +\log( \color{blue} b)\) \(\Large \log( a^{\color{green} n})= \color{green} n \log (a)\)

OpenStudy (anonymous):

so \[\log x +{1/2}\log x^2 +1\]

OpenStudy (anonymous):

Don't forget your parenthesis...it's incorrect otherwise.\[\log x + \frac 12 \log\color{red}( x^2 +1\color{red})\]